A student eats a dinner containing 2000Calories of energy. He wishes to do an equivalent amount of work in the gymnasium by lifting a 10Kg object. How many times must he raise the object to expand this much work? Assume that he lies to his back and raises it a distance of 0.5m each, and the effort done to let the weight down is 50% of the effort done to raise it.


1
Expert's answer
2021-11-09T10:58:46-0500

"E = 2000\ cal"

"E\approx 2000*4.2=8400\ J"

"m= 10kg"

"g = 9.8\frac{m}{s^2}"

"h =0.5\ m"

"\text{Energy for one lift in and one lowering of the weight:}"

"E_1= mgh +0.5mgh =1.5mgh"

"E_1= 1.5*10*9.8*0.5=73.5\ J"

"n =\lceil \frac{E}{E_1}\rceil= \lceil \frac{8400}{73.5}\rceil=115"


"\text{Answer: }115"